# An introduction to reinforcement learning for neuroscience

### v2

@@ 7 Deep reinforcement learning
We have so far considered small state and action spaces, where tabular policies are tractable. Unfortunately, most ethologically relevant state spaces are large enough that we cannot enumerate all possible states and actions. However, novel situations often resemble previously encountered states, allowing agents to generalize shared structure to these new but related settings (Botvinick et al., , 2020). In these cases, we can use function approximation (Sutton and Barto, , 2018) instead of tabular policies. This involves an assumption that similar states will have similar state-action values and should therefore have similar policies. By making this assumption, we can generalize to unseen states based on previous experience. The use of deep or recurrent neural networks as powerful function approximators for reinforcement learning has driven impressive progress in this setting – the domain of ‘deep reinforcement learning’ (deep RL). Deep RL has seen increasing interest not just in machine learning, but also as a model of neural dynamics and behaviour in humans and other animals (Wang et al., , 2018; Jensen et al., , 2023; Makino, , 2023; Merel et al., , 2019; Banino et al., , 2018; Aldarondo et al., , 2024; Botvinick et al., , 2020). Popular approaches in (model-free) deep RL can largely be divided into two categories: value-based methods, which compute state-action values that can be used for action selection; and policy gradient methods, which train a neural network to output a policy directly.
@@
The general recipe involves defining a state action value function $Q_{\theta}(s,a)$, where the parameters $\theta$ of the deep network defining our agent are learned as follows:
@@
  • Update the network parameters $\Delta\theta\propto-\frac{\partial\mathcal{L}}{\partial\theta}$.
@@
On the surface, this looks like a straightforward generalization of tabular Q-learning, and it may seem surprising that deep Q-learning did not see significant use or success until the foundational work of Mnih et al., (2013). However, a major difficulty arises from the autocorrelation of the states observed by the agent, which destabilizes training. This can be mitigated by the use of ‘experience replay’, where the experience generated by the agent is added to a global replay buffer $\mathcal{B}$. One or more experiences are then sampled randomly from the buffer at each iteration and used to update the network parameters – reminiscent of the ‘DYNA’ architecture described previously. An additional algorithmic instability arises from the term $\text{max}_{a}Q_{\theta}(s_{t+1},a)$, which we cannot differentiate despite its dependence on $\theta$. It is therefore common to use a ‘target network’ $Q_{\theta^{\prime}}(s_{t+1},a)$ that is fixed for a period of time, usually after being set to a copy of our ‘student network’. Together, these two approaches give rise to the ‘deep Q network’ (DQN) developed by Mnih et al., (2013), which is trained as follows:
@@
  • Define a loss $\mathcal{L}(\theta)=0.5[Q_{\theta}(s^{\prime}_{t},a^{\prime}_{t})-(r^{\prime}_{t}+\gamma\text{max}_{a}Q_{\theta^{\prime}}(s^{\prime}_{t+1},a))]^{2}$ [optionally averaged over the full batch]. Note the student network has parameters $\theta$ and the target network inside the max has parameters $\theta^{\prime}$.
@@
This algorithm is effectively off-policy, since most experience in $\mathcal{B}$ was collected by a policy defined by an old set of parameters – and the data in $\mathcal{B}$ can in fact be generated completely independently of the agent being trained. Even though the DQN is more stable than naive deep Q-learning, an additional instability arises from the fact that $Q_{\text{max}}(s^{\prime}_{t+1})=\text{max}_{a}Q_{\theta^{\prime}}(s^{\prime}_{t+1},a)$ uses the same Q values both to estimate which action is best and what the value of that action is, which leads to a positively biased estimate. This can be mitigated by ‘double Q-learning’ (Van Hasselt et al., , 2016), where the student network selects the best action and the target network evaluates its value, $Q_{\text{max}}(s^{\prime}_{t+1})\leftarrow Q_{\theta^{\prime}}(s^{\prime}_{t+1},\text{argmax}_{a}(Q_{\theta}(s^{\prime}_{t+1},a)))$.
While modern deep Q-learning has reached impressive performance across a range of machine learning settings (Mnih et al., , 2013; Lillicrap et al., , 2015; Schaul et al., , 2015; Kalashnikov et al., , 2018), it is unclear whether the various modifications needed to stabilize the algorithm could be implemented in biological circuits. This is perhaps the reason why neuroscience research using deep Q-learning has been relatively scarce, despite the prevalence of tabular Q-learning in the theoretical neuroscience literature. An interesting exception is recent work by Makino, (2023), which shows parallels between the values learned by a DQN and neural representations in mammalian cortex during a compositional behavioural task. Additionally, the importance of experience replay in DQNs (Mnih et al., , 2013; Schaul et al., , 2015) has close parallels to the proposal that hippocampal replay constitutes a form of experience replay (Mattar and Daw, , 2018).
@@
A conceptually simpler approach for deep reinforcement learning uses policy gradient methods (Sutton and Barto, , 2018), where a neural network with parameters $\theta$ takes as input the (observable) state of the environment and directly outputs a policy $\pi_{\theta}$. This has also found more support and use in the neuroscience literature, where policy gradient methods have recently been used as models of learning and neural dynamics in the biological brain (Wang et al., , 2018; Jensen et al., , 2023; Merel et al., , 2019; Song et al., , 2017).
@@
Inserting Equation 36 in Equation 33, we arrive at the REINFORCE algorithm (Williams, , 1992):
@@
This is the formulation most commonly used in the literature, but it is not actually the same as Equation 39, which would use $R_{t}=\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-0}r_{t^{\prime}}$. As briefly discussed in Section 2, this is because the discount factor $\gamma$ is generally used as a variance reduction method rather than because we intrinsically care less about rewards later in the task. In fact, Equation 40 is not strictly speaking a gradient (Nota and Thomas, , 2019), which is why we denote it $\hat{\nabla}$.
It can also be shown that subtracting an action-independent baseline from $R_{t}$ does not change the expectation in Equation 40, while potentially reducing its variance. A common choice is the expected future reward $V(s_{t})$, which gives rise to the so-called ‘actor-critic’ algorithm
@@
Intuitively, Equation 41 upregulates the probability of actions that lead to higher-than-expected reward and downregulates the probability of actions that lead to lower-than-expected reward. A rich neuroscience literature suggests a biological implementation of an actor-critic-like algorithm, with dorsal striatum implementing the ‘actor’ and ventral striatum the ‘critic’ (Takahashi et al., , 2008; Sutton and Barto, , 2018; O’Doherty et al., , 2004).
Finally, it is common to reduce the variance of the gradient estimate through an approach known as ‘bootstrapping’, which approximates $R_{t}\approx r_{t}+\gamma V(s_{t+1})$. This is useful because $r_{t}+\gamma V(s_{t+1})$ has lower variance than $\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}$. We therefore replace $(R_{t}-V(s_{t}))$ in Equation 41 with the ‘advantage function’ $A_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$ to implement the ‘advantage actor-critic’ (A2C) algorithm. In between these two extreme cases of a full Monte Carlo estimate of $R_{t}$ and a ‘one-step’ bootstrap, the sum in $R_{t}$ can be truncated to any order with $R_{t^{\prime}}$ replaced by $V(s_{t^{\prime}})$ (Sutton and Barto, , 2018). In theory, this gradient estimate remains unbiased if the value function is correct. In practice the estimate of $V(s_{t^{\prime}})$ learned by the critic will be inexact, which biases the parameter updates. Bootstrapping therefore leads to a tradeoff between the bias and variance of parameter updates.
@@
Figure 5: Meta-reinforcement learning. The results in this figure reproduce some of the analyses in Figure 1 of Wang et al., (2018). (A) We trained a recurrent meta-reinforcement learning agent in a two-armed bandit task, where the reward probabilities of each arm were sampled independently from $\mathcal{U}(0,1)$ at the beginning of each episode and remained fixed throughout the episode. A recurrent neural network was trained across many episodes with different reward probabilities using an actor-critic algorithm. The input to the agent consisted of the previous action, the previous reward, and the time-within-trial. The average reward per episode is plotted against the episode number, showing that the agent gradually learns to adapt within each episode to the particular instantiation of the bandit task. Importantly, the parameters of the network are fixed within an episode, meaning that this adaptation occurs through the recurrent dynamics. Dashed horizontal lines indicate the reward of an agent selecting random actions and an ‘oracle’ agent that always chooses the best arm. (B) Heatmap showing example behaviour of the agent in episodes with different reward probabilities for the first arm, $p(r|a=1)$. For the analysis here and in (C), we set $p(r|a=2)=1-p(r|a=1)$. Across episodes, the agent experiments with different actions and eventually converges on the optimal action. For episodes with more similar reward probabilities (near the middle), it takes longer to identify the optimal action. This balance between exploration and exploitation is mediated by the recurrent network dynamics, which are learned over many episodes using deep reinforcement learning. (C) We averaged the hidden state of the RNN over 100 episodes for each of several different reward probabilities, ranging from low (green) to high (blue) $p(r|a=1)$. We then performed PCA on the resulting matrix of average hidden states to compute a low-dimensional trajectory over the course of an episode for each reward probability. This two-dimensional embedding of neural activity converges to different regions of state space during the episode for different reward probabilities. Black cross indicates the hidden state at the beginning of an episode, and coloured points indicate the final hidden state in an episode for the different reward probabilities.
@@
While these policy gradient methods may seem far removed from neuroscience, they have had several applications in recent years (Wang et al., , 2018; Jensen et al., , 2023; Merel et al., , 2019; Li et al., , 2022; Song et al., , 2017). Of particular interest, Wang et al., (2018) suggested that frontal cortex can be modelled as a recurrent deep RL agent, where the RNN parameters are configured by learning from rewards over long periods of time from many tasks that have a shared underlying structure. Importantly, this ‘slow’ model-free learning process gives rise to an agent that can rapidly learn from experience with fixed parameters when exposed to a new task from the same distribution. This is achieved by the agent learning to effectively implement a fast RL-like algorithm in the dynamics of the network (Figure 5). This process, whereby an agent trained slowly on a large distribution of tasks can rapidly adapt to a new task, is known as ‘meta-reinforcement learning’ (Finn et al., , 2017; Ritter et al., , 2018; Duan et al., , 2016; Wang et al., , 2016). Wang et al., (2018) showed that such a meta-RL model can explain a range of neuroscientific findings, including
@@
  • Dynamic adaptation of the effective learning rate of an agent to the volatility of the environment (Behrens et al., , 2007).
@@
  • The emergence of ‘model-based’ behaviour in the ‘two-step’ task commonly used to distinguish between model-free and model-based RL (Miller et al., , 2017; Daw et al., , 2011).
@@
  • The ability of animals to get progressively faster at learning when exposed to multiple tasks with a consistent abstract task structure (Harlow, , 1949).
@@
Further experimental evidence for this meta-RL framework comes from Hattori et al., (2023), who showed that across-session learning in a reversal learning task relied on synaptic plasticity in orbitofrontal cortex (OFC; a subregion of PFC), while within-session learning relied on recurrent dynamics in OFC. Recently, Jensen et al., (2023) also extended the work of Wang et al., (2018) to allow the meta-RL network dynamics to update the policy from imagined experience using a learned model of the environment – reminiscent of DYNA, but now implemented in RNN dynamics instead of parameter updates.
@@ 8 Distributional reinforcement learning
In Equation 3 and Equation 12, we defined the expected future reward for a given state or state-action pair. The methods considered so far have only used such expectations as a learning signal. However, recent research suggests that animals may in fact estimate entire future reward distributions (Dabney et al., , 2020; Sousa et al., , 2023). These studies were inspired by findings that such distributional RL can improve the performance of artificial agents (Bellemare et al., , 2017, 2023; Dabney et al., , 2018). To formalize this, we use $Z^{\pi}(s,a)$ to denote a single sample from the distribution over possible cumulative discounted future rewards resulting from following policy $\pi$ after taking action $a$ in state $s$:
@@
Instead of only estimating this expectation, we now want to learn the full distribution of returns, $p_{z^{\pi}}(Z^{\pi}(s,a))$. One normative reason to learn this distribution is to develop methods that are risk averse (Morimura et al., , 2010, 2012) or explicitly take into account uncertainty (Dearden et al., , 1998). However, recent work has suggested that such a distributional approach can also increase expected reward by improving representation learning in the deep RL setting (Bellemare et al., , 2017; Dabney et al., , 2018; Rowland et al., , 2019; Bellemare et al., , 2023). This is because traditional deep RL only distinguishes states that have different expected value, while distributional RL learns to distinguish states that have different value distributions (Figure 6A).
@@
This is a generalization of the mean, which is recovered for $\tau=0.5$, similar to how the quantile generalizes the notion of a median. A distribution is uniquely specified by its expectiles, and we can therefore represent $p_{z^{\pi}}(Z^{\pi}(s,a))$ in terms of $\{\epsilon_{\tau}\}$. Translating this to an RL algorithm involves training a network (or tabular values) to predict a set of expectiles for a given state (and action). The parameters of the agent are then updated by propagating the distribution implied by the predicted expectiles through the Bellman equation and minimizing the difference between the initial and propagated distributions (Bellemare et al., , 2023).
In the tabular value learning case (c.f. Section 3), this turns out to be particularly simple to do using a modified TD-update rule (c.f. Equation 9; Lowet et al., , 2020; Figure 6B). In particular, we consider a set of units $\{V_{\tau_{i}}(s)\}$, each with a target expectile $\tau_{i}:=\frac{\alpha_{i}^{+}}{\alpha_{i}^{+}+\alpha_{i}^{-}}$ of the return distribution $p_{z^{\pi}}(Z^{\pi}(s))$ (where $p_{z^{\pi}}(Z^{\pi}(s))=\sum_{a}\left[\pi(a|s)p_{z^{\pi}}(Z^{\pi}(s))\right]$). These expectiles can be learned by sampling experience from the environment under the policy and defining a TD error for each unit and state transition as
@@
Here, $\tilde{Z}^{\pi}_{j}(s_{t+1})$ is a random sample from the learned approximate distribution of cumulative future returns from state $s_{t+1}$, $p_{\{V_{\tau_{i}}\}}(\tilde{Z}^{\pi}(s_{t+1}))$ (Lowet et al., , 2020; Dabney et al., , 2020). We then apply the following update rule to all units:
@@
Figure 6: Distributional reinforcement learning. (A) Example of distributional RL improving representation learning. Two states (e.g. ‘banana’ and ‘lemon’) may have the same expected reward but different reward distributions (top row). A standard RL agent only has to predict the mean (orange vertical lines) and may learn a simple predictive feature like ‘yellow’ that combines both states. If the mean reward now changes for one state (bottom row), e.g. because the agent learned to make a banana smoothie, it may erroneously generalize to all yellow fruits (bottom right; red line). However, a distributional agent is forced to learn an initial representation that distinguishes the states, which can improve downstream learning and prevent overgeneralization (blue curves). (B) Distributional RL simulations on a simple value estimation task with a single state and no actions, reproducing key ideas from Dabney et al., (2020). Distribution of rewards (black), plotted together with the learned values $V_{\tau_{i}}$ across 20 units learning through standard TD learning (orange; all $\tau_{i}=0.5$), or through distributional RL with different $\tau_{i}=\frac{\alpha_{i}^{+}}{\alpha_{i}^{+}+\alpha_{i}^{-}}$ (blue). The TD units all converge to the mean reward, while the expectile units end up tiling the distribution. (C) For both the TD units and distributional units from (B), we plot the temporal difference updates performed in response to different rewards from the reward distribution. These updates have been proposed to be represented in the firing rates of dopaminergic VTA neurons relative to baseline (Schultz et al., , 1997; Dabney et al., , 2020). The TD units show a constant linear scaling across positive and negative rewards, while the distributional units show an asymmetric scaling of firing rate with reward magnitude above and below their reversal point (black horizontal line). The ratio of slopes above and below the reversal point scales positively with the value of the reversal point. These features of dopaminergic VTA neurons were used by Dabney et al., (2020) to argue that the brain implements a form of distributional RL. (D) True reward distribution (black) reproduced from (B), now plotted together with the reward distribution $p_{\{V_{\tau_{i}}\}}(\tilde{Z}^{\pi}(s))$ implied by the distributional units from (B) and (C) at different stages of learning (green to blue). These imputed distributions were computed by assuming that the value $V_{\tau_{i}}$ learned by unit $i$ corresponds to expectile $\tau_{i}=\frac{\alpha_{i}^{+}}{\alpha_{i}^{+}+\alpha_{i}^{-}}$ of the reward distribution. We then infer the distribution implied by these expectiles under the assumption that it consists of a set of 20 delta functions (Rowland et al., , 2019). Finally, we convolved the resulting delta functions with a Gaussian kernel ($\sigma=0.1$) for visualization. This process was repeated using $\{V_{\tau_{i}}\}$ at different stages of learning. The units were all initialized at $V_{\tau_{i}}=0.5$, so the initial distribution is a delta function at $r=0.5$ (green). At the end of learning, the population faithfully represents the true reward distribution, capturing key features including bimodality and the relative magnitude of the two modes (blue). Dabney et al., (2020) used a similar approach to infer the distribution encoded by dopaminergic VTA neurons at the end of animal training and found a close match to the true reward distribution.
@@
Intriguingly, recent work in neuroscience has suggested that distributional RL could underlie value learning in biological neural circuits (Dabney et al., , 2020; Lowet et al., , 2020; Sousa et al., , 2023). In particular, Dabney et al., (2020) recorded the activity of dopaminergic VTA neurons during a task with stochastic rewards and showed that the neurons appeared to represent the full distribution of possible outcomes using an expectile representation. More concretely, they showed that:

### v3

@@ 7 Deep reinforcement learning
We have so far considered small state and action spaces, where tabular policies are tractable. Unfortunately, most ethologically relevant state spaces are large enough that we cannot enumerate all possible states and actions. However, novel situations often resemble previously encountered states, allowing agents to generalize shared structure to these new but related settings [14]. In these cases, we can use function approximation [108] instead of tabular policies. This involves an assumption that similar states will have similar state-action values and should therefore have similar policies. By making this assumption, we can generalize to unseen states based on previous experience. The use of deep or recurrent neural networks as powerful function approximators for reinforcement learning has driven impressive progress in this setting – the domain of ‘deep reinforcement learning’ (deep RL). Deep RL has seen increasing interest not just in machine learning, but also as a model of neural dynamics and behaviour in humans and other animals [123, 47, 62, 66, 6, 4, 14]. Popular approaches in (model-free) deep RL can largely be divided into two categories: value-based methods, which compute state-action values that can be used for action selection; and policy gradient methods, which train a neural network to output a policy directly.
@@
The simplest approach involves defining a state-action value function $Q_{\theta}(s,a)$, where the parameters $\theta$ of the neural network defining our agent are learned as follows:
@@
  • Update parameters $\Delta\theta\propto-\frac{\partial\mathcal{L}}{\partial\theta}$    [often treating the ‘target value’ $y_{t}=(r_{t}+\gamma\text{max}_{a}Q_{\theta}(s_{t+1},a))$ as constant w.r.t. $\theta$].
@@
On the surface, this looks like a straightforward generalization of tabular Q-learning, and it may seem surprising that deep Q-learning did not see significant use or success until the foundational work of Mnih et al. [71]. However, a major difficulty arises from the autocorrelation of the states observed by the agent, which destabilizes training. This can be mitigated by the use of ‘experience replay’, where the experience generated by the agent is added to a global replay buffer $\mathcal{B}$. One or more experiences are then sampled randomly from the buffer at each iteration and used to update the network parameters – reminiscent of the ‘Dyna’ architecture described previously. Additionally, the target value $y_{t}=(r_{t}+\gamma\text{max}_{a}Q_{\theta}(s_{t+1},a))$ itself depends on $\theta$ and therefore changes when any Q-value is updated (in contrast to tabular Q-learning, where there is no parameter sharing). It is therefore common to use a ‘target network’ $Q_{\theta^{\prime}}$ that remains fixed for multiple rounds of data collection and parameter updates. This reduces fluctuations in the target values, and the resulting parameter updates are gradients of a well-defined objective function. Together, these two approaches give rise to the ‘deep Q network’ (DQN) developed by Mnih et al. [71], which is trained as follows:
@@
  • Define a loss $\mathcal{L}(\theta)=0.5[Q_{\theta}(s^{\prime}_{t},a^{\prime}_{t})-(r^{\prime}_{t}+\gamma\text{max}_{a}Q_{\theta^{\prime}}(s^{\prime}_{t+1},a))]^{2}$    [optionally averaged over the full batch]. Note the ‘student network’ has parameters $\theta$ and the target network inside the max has parameters $\theta^{\prime}$.
@@
This algorithm is effectively off-policy, since most of the data in $\mathcal{B}$ is collected by a policy defined by an old set of parameters – and the data in $\mathcal{B}$ can in fact be generated completely independently of the agent being trained. Even though the DQN is more stable than naive deep Q-learning, an additional instability arises from the fact that $Q_{\text{max}}(s^{\prime}_{t+1})=\text{max}_{a}Q_{\theta^{\prime}}(s^{\prime}_{t+1},a)$ uses the same Q values both to estimate which action is best and what the value of that action is, which leads to a positively biased estimate. This can be mitigated by ‘double Q-learning’ [118], where the student network selects the best action and the target network evaluates its value, $Q_{\text{max}}(s^{\prime}_{t+1})\leftarrow Q_{\theta^{\prime}}(s^{\prime}_{t+1},\text{argmax}_{a}(Q_{\theta}(s^{\prime}_{t+1},a)))$.
While modern deep Q-learning has reached impressive performance across a range of machine learning settings [71, 59, 94, 51], it is unclear whether the various modifications needed to stabilize the algorithm could be implemented in biological circuits. This is perhaps the reason why neuroscience research using deep Q-learning has been relatively scarce, despite the prevalence of tabular Q-learning in theoretical neuroscience. An interesting exception is recent work by Makino [62], which shows parallels between the values learned by a DQN and neural representations in mammalian cortex during a compositional behavioural task. Additionally, the importance of experience replay in DQNs [71, 94] has close parallels to the proposal that hippocampal replay constitutes a form of experience replay [64].
@@
A conceptually simpler approach for deep reinforcement learning uses policy gradient methods [108], where a neural network with parameters $\theta$ takes as input the (observable) state of the environment and directly outputs a policy $\pi_{\theta}$. This has also found more support and use in the neuroscience literature, where policy gradient methods have recently been used as models of learning and neural dynamics in the biological brain [123, 47, 66, 101].
@@
Inserting Equation 34 in Equation 31, we arrive at the REINFORCE algorithm [130]:
@@
This is the formulation most commonly used in the literature, but it is not actually the same as Equation 37, which would use $R_{t}=\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-0}r_{t^{\prime}}$. As briefly discussed in Section 2, this is because the discount factor $\gamma$ is generally used as a variance reduction method rather than because we intrinsically care less about rewards later in the task. In fact, Equation 38 is not strictly speaking a gradient [77], which is why we denote it $\hat{\nabla}$.
It can also be shown that subtracting an action-independent baseline from $R_{t}$ does not change the expectation in Equation 38, while potentially reducing its variance. A common choice is the expected future reward $V(s_{t})$:
@@
Intuitively, Equation 39 upregulates the probability of actions that lead to higher-than-expected reward and downregulates the probability of actions that lead to lower-than-expected reward.
Finally, it is common to reduce the variance of the gradient estimate further through an approach known as ‘bootstrapping’, which approximates $R_{t}\approx r_{t}+\gamma V(s_{t+1})$. This is useful because $r_{t}+\gamma V(s_{t+1})$ has lower variance than $\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}$. We therefore replace $(R_{t}-V(s_{t}))$ in Equation 39 with the ‘advantage function’ $A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})\approx r_{t}+\gamma V(s_{t+1})-V(s_{t})$. In between these two extreme cases of a full Monte Carlo estimate of $R_{t}$ and a ‘one-step’ bootstrap, the sum in $R_{t}$ can be truncated to any order, with $R_{t^{\prime}}$ replaced by $V(s_{t^{\prime}})$ [108]. In theory, this gradient estimate remains unbiased if the value function is correct. In practice, the learned estimate of $V(s_{t^{\prime}})$ will be inexact, which biases the parameter updates. Bootstrapping therefore leads to a tradeoff between the bias and variance of parameter updates.
These variance reduction approaches give rise to the so-called ‘actor-critic’ algorithm, where an agent both computes a policy $\pi$ (the actor) and an ‘evaluation’ of the policy in the form of state(-action) values (the critic). A rich neuroscience literature suggests that the basal ganglia of biological agents implement an actor-critic-like algorithm. Here, dorsal striatum is proposed to implement the ‘actor’ and ventral striatum the ‘critic’ [110, 108, 79].
@@
While these policy gradient methods may seem far removed from neuroscience, it has been found that neural networks trained with policy gradients often learn representations and behaviours reminiscent of biological organisms [123, 47, 66, 58, 101].
Meta-reinforcement learning
@@
Figure 5: Meta-reinforcement learning. The results in this figure reproduce some of the analyses in Figure 1 of Wang et al. [123]. (A) We trained a recurrent meta-reinforcement learning agent in a two-armed bandit task, where the reward probabilities of each arm were sampled independently from $\mathcal{U}(0,1)$ at the beginning of each episode and remained fixed throughout the episode. A recurrent neural network was trained across many episodes with different reward probabilities using an actor-critic algorithm. The input to the agent consisted of the previous action, the previous reward, and the time-within-trial. The average reward per episode is plotted against the episode number, showing that the agent gradually learns to adapt within each episode to the particular instantiation of the bandit task. Importantly, the parameters of the network are fixed within an episode, meaning that this adaptation occurs through the recurrent dynamics. Dashed horizontal lines indicate the reward of an agent selecting random actions and an ‘oracle’ agent that always chooses the best arm. (B) Heatmap showing example behaviour of the agent in episodes with different reward probabilities for the first arm, $p(r|a=1)$. For the analysis here and in (C), we set $p(r|a=2)=1-p(r|a=1)$. Across episodes, the agent experiments with different actions and eventually converges on the optimal action. For episodes with more similar reward probabilities (near the middle), it takes longer to identify the optimal action. This balance between exploration and exploitation is mediated by the recurrent network dynamics, which are learned over many episodes using deep reinforcement learning. (C) We averaged the hidden state of the RNN over 100 episodes for each of several different reward probabilities, ranging from low (green) to high (blue) $p(r|a=1)$. We then performed PCA on the resulting matrix of average hidden states to compute a low-dimensional trajectory over the course of an episode for each reward probability. This two-dimensional embedding of neural activity converges to different regions of state space during the episode for different reward probabilities. Black cross indicates the hidden state at the beginning of an episode, and coloured points indicate the final hidden state in an episode for the different reward probabilities.
@@
A prominent example of deep reinforcement learning providing insights into biological circuits is the ‘recurrent meta-reinforcement learning’ model developed by Wang et al. [123]. The authors trained a recurrent deep RL agent using policy gradients, where the RNN parameters are configured by learning from rewards over long periods of time from many tasks that have a shared underlying structure. Importantly, this ‘slow’ model-free learning process gives rise to an agent that can rapidly learn from experience with fixed parameters when exposed to a new task from the same task distribution. This is achieved by the agent learning to effectively implement a fast RL-like algorithm in the dynamics of the network (Figure 5). This process, whereby an agent trained slowly on a large distribution of tasks can rapidly adapt to a new task, is known as ‘meta-reinforcement learning’ [34, 90, 28, 124]. Wang et al. [123] suggested that prefrontal cortex resembles such a recurrent meta-RL system, and their model explained a range of neuroscientific findings. This included
@@
  • Dynamic adaptation of the effective learning rate of an agent to the volatility of the environment [8].
@@
  • The emergence of ‘model-based’ behaviour in the ‘two-step’ task commonly used to distinguish between model-free and model-based RL [68, 22].
@@
  • The ability of animals to get progressively faster at learning when exposed to multiple tasks with a consistent abstract task structure [41].
@@
Further experimental evidence for this meta-RL framework comes from Hattori et al. [42], who showed that across-session learning in a reversal learning task relied on synaptic plasticity in orbitofrontal cortex (OFC; a subregion of PFC), while within-session learning relied on recurrent dynamics in OFC. Recently, Jensen et al. [47] also extended the work of Wang et al. [123] to allow the meta-RL network dynamics to update the policy from imagined experience using a learned model of the environment – reminiscent of Dyna, but now implemented in RNN dynamics instead of parameter updates.
@@ 8 Distributional reinforcement learning
In Equation 3 and Equation 10, we defined the expected future reward for a given state or state-action pair. The methods considered so far have only used such expectations as a learning signal. However, recent research suggests that animals may in fact estimate entire future reward distributions [20, 102]. These studies were inspired by findings that such distributional RL can improve the performance of artificial agents [9, 10, 21]. To formalize this, we use $Z^{\pi}(s,a)$ to denote a single sample from the distribution over possible cumulative discounted future rewards resulting from following policy $\pi$ after taking action $a$ in state $s$:
@@
Instead of only estimating this expectation, we now want to learn the full distribution of returns, $p_{z^{\pi}}(Z^{\pi}(s,a))$. One normative reason to learn this distribution is to develop methods that are risk averse [74, 73] or explicitly take into account uncertainty [25]. However, recent work has suggested that such a distributional approach can also increase expected reward by improving representation learning in the deep RL setting [9, 21, 92, 10]. This is because traditional deep RL only distinguishes states that have different expected value, while distributional RL learns to distinguish states that have different value distributions (Figure 6A).
@@
This is a generalization of the mean, which is recovered for $\tau=0.5$, similar to how the quantile generalizes the notion of a median. A distribution is uniquely specified by its expectiles, and we can therefore represent $p_{z^{\pi}}(Z^{\pi}(s,a))$ in terms of $\{\epsilon_{\tau}\}$. Translating this to an RL algorithm involves training a network (or tabular values) to predict a set of expectiles for a given state (and action). The parameters of the agent are then updated by propagating the distribution implied by the predicted expectiles through the Bellman equation and minimizing the difference between the initial and propagated distributions [10].
In the tabular value learning case (c.f. Section 3), this can be achieved using a modified TD-update rule (c.f. Equation 9; 61; Figure 6B). In particular, we consider a set of units $\{V_{\tau_{i}}(s)\}$, each with a target expectile $\tau_{i}:=\frac{\alpha_{i}^{+}}{\alpha_{i}^{+}+\alpha_{i}^{-}}$ of the return distribution $p_{z^{\pi}}(Z^{\pi}(s))$ (where $p_{z^{\pi}}(Z^{\pi}(s))=\sum_{a}\left[\pi(a|s)p_{z^{\pi}}(Z^{\pi}(s))\right]$). These expectiles can be learned by sampling experience from the environment under the policy and defining a TD error for each unit and state transition as
@@
Here, $\tilde{Z}^{\pi}_{j}(s_{t+1})$ is a random sample from the learned approximate distribution of cumulative future returns from state $s_{t+1}$, $p_{\{V_{\tau_{i}}\}}(\tilde{Z}^{\pi}(s_{t+1}))$ [61, 20]. We then apply the following update rule to all units:
@@
Figure 6: Distributional reinforcement learning. (A) Example of distributional RL improving representation learning. Two states (e.g. ‘banana’ and ‘lemon’) may have the same expected future reward but different reward distributions (top row). A standard RL agent only has to predict the mean (orange vertical lines) and may learn a simple predictive feature like ‘yellow’ that combines both states. If the expected future reward now changes for one state (bottom row), e.g. because the agent learned to make a banana smoothie, it may erroneously generalize to all yellow fruits (bottom right; red line). However, a distributional agent is forced to learn an initial representation that distinguishes the states, which can improve downstream learning and prevent overgeneralization (blue curves). (B) Distributional RL simulations on a value estimation task with a single state and no actions, reproducing key ideas from Dabney et al. [20]. Distribution of rewards (black), plotted together with the learned values $V_{\tau_{i}}$ across 20 units learning through standard TD learning (orange; all $\tau_{i}=0.5$), or through distributional RL with different $\tau_{i}=\frac{\alpha_{i}^{+}}{\alpha_{i}^{+}+\alpha_{i}^{-}}$ (blue). The TD units all converge to the mean reward, while the expectile units end up tiling the distribution. (C) For both the TD units and distributional units from (B), we plot the temporal difference updates performed in response to different rewards from the reward distribution. These updates have been proposed to be represented in the firing rates of dopaminergic VTA neurons relative to baseline [97, 20]. The TD units show a constant linear scaling across positive and negative rewards, while the distributional units show an asymmetric scaling of firing rate with reward magnitude above and below their reversal point (black horizontal line). The ratio of slopes above and below the reversal point scales positively with the value of the reversal point. These features of dopaminergic VTA neurons were used by Dabney et al. [20] to argue that the brain implements a form of distributional RL. (D) True reward distribution (black) reproduced from (B), now plotted together with the reward distribution $p_{\{V_{\tau_{i}}\}}(\tilde{Z}^{\pi}(s))$ implied by the distributional units from (B) and (C) at different stages of learning (green to blue). These imputed distributions were computed by assuming that the value $V_{\tau_{i}}$ learned by unit $i$ corresponds to expectile $\tau_{i}=\frac{\alpha_{i}^{+}}{\alpha_{i}^{+}+\alpha_{i}^{-}}$ of the reward distribution. We then infer the distribution implied by these expectiles under the assumption that it consists of a set of 20 delta functions [92]. Finally, we convolved the resulting delta functions with a Gaussian kernel ($\sigma=0.1$) for visualization. This process was repeated using $\{V_{\tau_{i}}\}$ at different stages of learning. The units were all initialized at $V_{\tau_{i}}=0.5$, so the initial distribution is a delta function at $r=0.5$ (green). At the end of learning, the population faithfully represents the true reward distribution, capturing key features including bimodality and the relative magnitude of the two modes (blue). Dabney et al. [20] used a similar approach to infer the distribution encoded by dopaminergic VTA neurons at the end of animal training and found a close match to the true reward distribution.
@@
Intriguingly, recent work in neuroscience suggests that distributional RL could underlie value learning in biological neural circuits [20, 61, 102]. In particular, Dabney et al. [20] recorded the activity of dopaminergic VTA neurons during a task with stochastic rewards and showed that the neurons appeared to represent the full distribution of possible outcomes using an expectile representation. More concretely, they showed that: